# Active Sensing for Multiuser Beam Tracking with Reconfigurable Intelligent Surface

### v1

@@ I-A Related Works
The problem setting in this paper also requires the modeling of spatial relationship between the AP, the RIS, and the UEs. Toward this end, GNN is an indispensable tool for learning these spatial relations [20, 21, 22]. Specifically, previous studies including [23, 20, 18] propose using GNNs to efficiently manage the mutual interference among the UEs by matching the graph structure of the neural network to the spatial connections between the UEs and the RIS. This is an important architectural feature that we also adopt in this paper. It allows more efficient training and gives more interpretable results.
One of the key differentiating features of this paper as compared to prior works is in the design of active sensing strategies for channel sensing. In most prior works [9, 10, 11, 12, 13, 15, 16, 17], the CSI acquisition stage uses fixed RIS reflection coefficients, either generated randomly or based on fixed patterns such as the discrete Fourier transform (DFT) [24]. However, these non-adaptive RIS sensing schemes are far from optimal. On the other hand, [18] adaptively designs the RIS sensing scheme based on an RNN, but proposes to employ the same RIS reflection coefficients in both sensing and data transmission stages. This is suboptimal due to the different functionalities of the RIS in two stages. A main contribution of this paper is a learning approach for adaptively designing the RIS coefficients in the channel sensing stage for tracking multiple UEs, which to the best of authors’ knowledge has not been explored in the existing literature. Recently, active sensing has demonstrated remarkable performance in related applications including beam alignment [25, 26, 6, 27, 28] and localization [29] in fixed channel environment, where the goal is to leverage previously obtained channel measurements to gradually refine focusing onto some desired low-dimensional part of the static channel. The current investigation is concerned with active sensing for beam tracking in the time-varying channels.
@@ I-B1 Active Sensing using RNN
This paper proposes to leverage historical channel information to adaptively design RIS reflection coefficients across multiple pilot stages to address a multiuser beam tracking problem. This is accomplished by adopting an LSTM-based RNN for its ability to capture the temporal correlations in the channel dynamics. At each CSI acquisition stage, a set of LSTM cells receive a new round of pilots from the mobile UEs, then update its fixed-dimensional hidden and cell state vectors accordingly. The cell state vectors are subsequently utilized to design the RIS reflection coefficients for the next round of channel sensing (as well as for data transmission). The proposed active sensing strategy can achieve significant performance gain over the benchmarks with non-adaptive sensing schemes.
@@ I-B2 Reflection Coefficients Optimization using GNN
This paper proposes to use GNN to learn the mappings from the LSTM cell state vectors to the RIS reflection coefficients for both channel sensing and data transmission. The GNN leverages the spatial relation between the UEs and the RIS to design the reflection patterns while accounting for the mutual interference between the UE’s. The proposed GNN architecture extends the designs in [18, 23, 30], while specifically addressing the distinct roles of the RIS reflection patterns in the sensing vs. data transmission stages. The proposed GNN improves upon previous work [18] that uses the same RIS coefficients for both sensing and communication. The proposed design also exhibits generalizability across different numbers of mobile UEs.
@@ IV-A GNN for Interference Management
To capture the spatial structure between the RIS and UEs, the initial graph representation vector $\bm{s}^{(t)}_{\xi,0}$’s undergo $D$ iterations of updates, each utilizing the output of the preceding layer as input. More specifically, the updates for the RIS node $w$ and $V$ incorporate their current representations as well as the representations of all UE nodes, while the representation vector of the UE node $k$ is updated based on its own prior representation and the representations of all other nodes in the network. These update procedures maintain the GNN’s permutation invariance and equivariance, and therefore provides better scalability and generalizability. The updating rules of the $d$-th updating layer are given as:
@@
After $D$ layers of iteration, the updated representation vector $\bm{s}^{(t)}_{\xi,D}$’s would contain sufficient information to design the beam tracking matrices. To obtain the correct output dimensions, we pass $\bm{s}^{(t)}_{k,D}$’s, $\bm{s}^{(t)}_{w,D}$, and $\bm{s}^{(t)}_{V,D}$ through linear layers $g_{1}(\cdot),g_{2}(\cdot),g_{3}(\cdot)$ with $2K$, $2N_{r}$, and $2N_{r}L$ fully connected units, respectively. Specifically, $\bm{s}^{(t)}_{k,D+1}=g_{1}(\bm{s}^{(t)}_{k,D})\in\mathbb{R}^{2K}$, $\bm{s}^{(t)}_{w,D+1}=g_{2}(\bm{s}^{(t)}_{w,D})\in\mathbb{R}^{2N_{r}}$, and $\bm{s}^{(t)}_{V,D+1}=g_{3}(\bm{s}^{(t)}_{V,D})\in\mathbb{R}^{2N_{r}L}$. The final representation vector $\bm{s}^{(t)}_{\xi,D+1}$’s are then normalized to produce the RIS sensing vectors $\{\bm{v}^{(t+1)}_{\ell}\}^{L}_{\ell=1}$, downlink RIS reflection coefficients $\bm{w}^{(t)}$, and power allocations $\{\bm{p}^{(t)},\bm{\lambda}^{(t)}\}$, respectively, while ensuring that the unit modulus constraints on $\bm{w}^{(t)}$, $\bm{v}_{\ell}^{(t+1)}$’s and the power constraints on $\bm{p}^{(t)},\bm{\lambda}^{(t)}$ are satisfied.
@@ V-A Simulation Setup
We consider an RIS-assisted MU-MISO communication system as illustrated in Fig. 8, consisting of an AP with $8$ antennas and an RIS with $100$ passive elements. In the Cartesian coordinate system shown in Fig. 8, the $(x,y,z)$-coordinates of the AP and the RIS are $(100\text{m},-100\text{m},0\text{m})$ and $(0\text{m},0\text{m},0\text{m})$, respectively. We assume that the RIS is equipped with a $10\times 10$ uniform rectangular array placed on the $(y,z)$-plane, and the AP has a uniform linear array parallel to the $x$-axis. The system is assumed to operate at a carrier frequency $f_{c}=1\text{GHz}$ with $10$MHz bandwidth. Unless stated otherwise, the uplink pilot transmit power $P_{u}$ and the downlink data transmit power $P_{d}$ are set to be $5$dBm and $10$dBm, and the noise spectrum density at the AP and UE are set to be $N_{0,u}=-140$dBm/Hz and $N_{0,d}=-160$dBm/Hz, respectively.
@@ V-B Implementation Details
The proposed active sensing framework is implemented on Tensorflow [39]. For the LSTM cells, the dimensions of hidden state $\bm{h}_{k}^{(t)}$’s, cell state $\bm{c}_{k}^{(t)}$’s, and hidden layers $\bm{u}_{c}$, $\bm{u}_{q}$, $\bm{u}_{i}$, $\bm{u}_{o}$, $\bm{w}_{c}$, $\bm{w}_{q}$, $\bm{w}_{i}$, $\bm{w}_{o}$ are set to be $512$. We adopt a GNN with $D=2$ updating layers. Within the GNN architecture, the DNNs $f_{1}(\cdot)$, $f_{2}(\cdot)$, and $f_{3}(\cdot)$ are of size $512\times 1024\times 512$, and the size of DNNs $f_{4}(\cdot)$, $f_{5}(\cdot)$, $f_{6}(\cdot)$, $f_{7}(\cdot)$, $f_{8}(\cdot)$, $f_{9}(\cdot)$, $f_{w}(\cdot)$, $f_{V}(\cdot)$ are $512\times 512\times 512$. We assume there are $N_{G}=8$ paths between the AP and RIS. Each transmission frame contains $N=3$ blocks. In the pilot stage (the 00-th block) of each frame, there are in total $4$ sub-blocks, i.e., $L=3$. The effective downlink channel $\bm{h}_{c,k}^{(t,0)}$’s are estimated in the $4$-th sub-block with $\tau_{w}=1000$ pilots. We concatenate $U=20$ active sensing units to train the proposed framework. In each training epoch, the neural network samples $6400$ training data including random channel and noise realizations. We terminate the training process if the validation loss does not decrease over $30$ consecutive training epochs. An instance of trajectories for $K=3$ mobile UEs is shown in Fig. 8.
@@ V-C Benchmark Methods
(a) Performance comparison with $L=3$.
(b) Generalizability across different UE numbers $K$.
(c) Impact of different $L$’s.
Figure 9: Average minimum user rate at the $1$-st block of each frame (i.e., $\min_{k}\{R_{k}^{(t,1)}\}$) for different beam tracking methods. Here, $M=8$, $N_{r}=100$, $N_{G}=8$, $P_{u}=5\text{dBm}$, $P_{d}=10\text{dBm}$. In (a), the neural networks are trained and tested with $L=3$ and $K=3$. In (b), neural networks are trained with $L=3$ and $K=3$, but tested under differnt $K$’s. In (c), the neural networks are trained and tested with $K=3$ and $L=1,3,10$.
@@ V-D Performance Comparison
Fig. 9(a) shows the downlink minimum user rate of different beam tracking methods in different transmission frames. The proposed active sensing framework initially achieves around $80\%$ of the BCD algorithm with perfect CSI, indicating efficient initial design based on the statistics of the UE mobility pattern. As the AP receives more pilots, the performance of the proposed approach improves quickly to $97\%$ of the performance of the benchmark with perfect CSI. This indicates that the proposed active sensing framework can efficiently exploit the temporal correlations of the time varying channel. Specifically, the beam tracking methods using both GNN and LSTM significantly outperform those benchmarks with only GNNs. This is because the LSTM-based methods are able to exploit the historical channel information to design the AP beamforming vectors and the downlink RIS reflection coefficients, thereby improving beam tracking performance over time. Among the beam tracking approaches with LSTM, the proposed framework performs the best, because it designs the RIS sensing vectors in an active manner. This shows the benefits of adaptively designing the RIS sensing vectors based on previously received pilots.
For the benchmarks with fixed sensing schemes, Fig. 9(a) shows that the ones with learned sensing vectors outperform those with random sensing vectors. This shows the performance gain of designing the RIS sensing vectors based on channel statistics. Further, we observe that the GNN & LSTM benchmark, which uses the same RIS reflection coefficients for both pilot acquisition and data transmission stages, performs less effectively as compared to other GNN & LSTM benchmarks that employ distinct RIS reflection coefficients for these two stages. This shows the importance of designing the RIS reflection coefficients in channel sensing and data transmission stages separately. Finally, we remark that a $10\%$ performance improvement can be obtained from power allocation refinement.
Generalizability
Next, we investigate the generalization capability of the proposed framework across different numbers of the mobile UEs $K$. As shown in Fig. 9(b), the proposed framework trained with $K=3$ is able to generalize to different numbers of UEs, and it always achieves more than $85\%$ performance of the perfect CSI benchmark. Additionally, the proposed active sensing framework can consistently improve its tracking performance by more than $15\%$ compare to the GNN benchmark with fixed sensing scheme.
@@
Moreover, we evaluate the performance of the proposed approach for different number of sensing vector $L$’s within each frame. In simulations, the noise spectrum density at the AP is set to $N_{0,u}=-135$dBm/Hz. As can be seen from Fig. 9(c), the proposed active sensing framework always outperforms the benchmarks with fixed sensing strategies, because it leverages past observations to actively design the RIS sensing vectors. Furthermore, it is observed that LSTM-based methods (including the proposed approach) show faster convergence as the number of sensing sub-blocks $L$ increases. This is because a larger $L$ provides the LSTM with more observations in every frame, thereby facilitating more efficient capturing of temporal channel correlations.
@@ V-E Interpretation of the Learned Solutions
(a) $\bm{v}^{(t)}_{1}$, actively designed by the proposed approach.
(b) Fixed $\bm{v}_{1}$, learned from channel statistics.
(c) $\bm{v}^{(t)}_{2}$, actively designed by the proposed approach.
(d) Fixed $\bm{v}_{2}$, learned from channel statistics.
(e) $\bm{v}^{(t)}_{3}$, actively designed by the proposed approach.
(f) Fixed $\bm{v}_{3}$, learned from channel statistics.
Figure 10: In the pilot stage (i.e., 00-th block) across different transmission frames, the array response of the RIS sensing vectors obtained by the proposed framework and the GNN & LSTM benchmark without active sensing [6]. Here, $L=3$, $M=8$, $N_{r}=100$, $K=3$, $N_{G}=8$, $P_{u}=5\text{dBm}$, $P_{d}=10\text{dBm}$.
@@
As illustrated in Fig. 10, the proposed framework designs $L=3$ RIS sensing vectors cooperatively. For example, in the pilot stage of the $36$-th frame, one UE (represented by the red dot) is not targeted by the first RIS sensing vector $\bm{v}_{1}^{(t)}$, but is focused by $\bm{v}_{2}^{(t)}$. Meanwhile, the weak array response of $\bm{v}_{2}^{(t)}$ towards the other two UEs is compensated by a stronger response from $\bm{v}_{3}^{(t)}$. This cooperative sensing strategy is not seen in the initial tracking frame, but is progressively developed over time by leveraging historical channel observations. As a result, the proposed framework actively narrows the sensing range and enhances the pilot SNR in the uplink CSI acquisition stages, leading to more informative neural network inputs and improved beam tracking performance. In contrast, the GNN & LSTM benchmark employs a non-adaptive approach by using channel statistics to train fixed RIS sensing vectors, which cover nearly the entire service area for CSI acquisition. However, this broad coverage results in lower uplink SNR during CSI acquisition and, subsequently, higher pilot overhead.
(a) Array response of the downlink RIS reflection coefficients obtained by the proposed framework.
(b) Array response of the downlink beamforming vectors obtained by the proposed framework.
Figure 11: Array response of the downlink RIS reflection coefficients and beamformers obtained by the proposed framework in different transmission frames. The proposed framework is tested on a randomly generated moving instance, with $L=3$, $M=8$, $N_{r}=100$, $K=3$, $N_{G}=8$, $P_{u}=5\text{dBm}$, $P_{d}=10\text{dBm}$.
Interpretation of the Downlink Beam Alignment Strategy
In Fig. 11(a), the array response of the downlink RIS reflection coefficients $\bm{w}^{(t)}$ shows that the proposed framework effectively tracks the UEs by progressively reflecting increasingly focused downlink beams via the RIS, towards the actual positions of the mobile UEs. It is observed that the peaks corresponding to the first two UEs are sometimes weaker than the third UE, but we can see from Fig. 11(b) that these weaker UEs are compensated by stronger array response of the beamformers at the AP. Therefore, the proposed framework indeed learns to jointly optimize the downlink RIS reflection coefficients and the beamforming matrix to ensure fairness in maximizing the minimum rate across the three mobile UEs, while exploiting channel correlations to design improved RIS reflection coefficients and beamformers over time.

### v2

@@ I-A Related Works
The problem setting in this paper also requires the modeling of spatial relationship between the AP, the RIS, and the UEs. Toward this end, GNN is an indispensable tool for learning these spatial relations [20, 21, 22, 23]. Specifically, previous studies including [24, 20, 23, 18] propose using GNNs to efficiently manage the mutual interference among the UEs by matching the graph structure of the neural network to the spatial connections between the UEs and the RIS. This is an important architectural feature that we also adopt in this paper. It allows more efficient training and gives more interpretable results.
One of the key differentiating features of this paper as compared to prior works is in the design of active sensing strategies for channel sensing. In most prior works [9, 10, 11, 12, 13, 15, 16, 17], the CSI acquisition stage uses fixed RIS reflection coefficients, either generated randomly or based on fixed patterns such as the discrete Fourier transform (DFT) [25]. However, these non-adaptive RIS sensing schemes are far from optimal. On the other hand, [18] adaptively designs the RIS sensing scheme based on an RNN, but proposes to employ the same RIS reflection coefficients in both sensing and data transmission stages. This is suboptimal due to the different functionalities of the RIS in two stages. A main contribution of this paper is a learning approach for adaptively designing the RIS coefficients in the channel sensing stage for tracking multiple UEs, which to the best of authors’ knowledge has not been explored in the existing literature. Recently, active sensing has demonstrated remarkable performance in related applications including localization [26] and beam alignment [27, 28, 6, 29, 30] in fixed channel environment, where the goal is to leverage previously obtained channel measurements to gradually refine focusing onto some desired low-dimensional part of the static channel. The current investigation is concerned with active sensing for beam tracking in the time-varying channels.
@@ I-B1 Active Sensing using RNN
This paper proposes to leverage historical channel information to adaptively design RIS reflection coefficients across multiple pilot stages to address a multiuser beam tracking problem. This is accomplished by adopting an LSTM-based RNN for its ability to capture the temporal correlations in the channel dynamics. At each CSI acquisition stage, a set of LSTM cells receive a new round of pilots from the mobile UEs, then update its fixed-dimensional hidden and cell state vectors accordingly. The cell state vectors are subsequently utilized to design the RIS reflection coefficients for the next round of channel sensing (as well as for data transmission). The proposed active sensing strategy can achieve significant performance gain over the benchmarks with non-adaptive sensing schemes. The proposed design also exhibits robust performance beyond the training range as it is able to update channel information based on the latest received pilots.
@@ I-B2 Reflection Coefficients Optimization using GNN
This paper proposes to use GNN to learn the mappings from the LSTM cell state vectors to the RIS reflection coefficients for both channel sensing and data transmission. The GNN leverages the spatial relation between the UEs and the RIS to design the reflection patterns while accounting for the mutual interference between the UE’s. The proposed GNN architecture extends the designs in [18, 24, 31], while specifically addressing the distinct roles of the RIS reflection patterns in the sensing vs. data transmission stages. The proposed GNN improves upon previous work [18] that uses the same RIS coefficients for both sensing and communication.
@@ IV-A GNN for Interference Management
To capture the spatial structure between the RIS and UEs, the initial graph representation vector $\bm{s}^{(t)}_{\xi,0}$’s undergo $D$ iterations of updates, each utilizing the output of the preceding layer as input. More specifically, the updates for the RIS node $w$ and $V$ incorporate their current representations as well as the representations of all UE nodes, while the representation vector of the UE node $k$ is updated based on its own prior representation and the representations of all other nodes in the network. These update procedures maintain the GNN’s permutation invariance and equivariance, and therefore allows more efficient training. The updating rules of the $d$-th updating layer are given as:
@@
After $D$ layers of iteration, the updated representation vector $\bm{s}^{(t)}_{\xi,D}$’s would contain sufficient information to design the beam tracking matrices. To obtain the correct output dimensions, we pass $\bm{s}^{(t)}_{k,D}$’s, $\bm{s}^{(t)}_{w,D}$, and $\bm{s}^{(t)}_{V,D}$ through linear layers $g_{1}(\cdot),g_{2}(\cdot),g_{3}(\cdot)$ with $2K$, $2N_{r}$, and $2N_{r}L$ fully connected units, respectively. Specifically, $\bm{s}^{(t)}_{k,D+1}=g_{1}(\bm{s}^{(t)}_{k,D})\in\mathbb{R}^{2K}$, $\bm{s}^{(t)}_{w,D+1}=g_{2}(\bm{s}^{(t)}_{w,D})\in\mathbb{R}^{2N_{r}}$, and $\bm{s}^{(t)}_{V,D+1}=g_{3}(\bm{s}^{(t)}_{V,D})\in\mathbb{R}^{2N_{r}L}$. The final representation vector $\bm{s}^{(t)}_{\xi,D+1}$’s are then normalized to produce the RIS sensing vectors $\{\bm{v}^{(t+1)}_{\ell}\}^{L}_{\ell=1}$, downlink RIS reflection coefficients $\bm{w}^{(t)}$, and power allocations $\{\bm{p}^{(t)},\bm{\lambda}^{(t)}\}$, respectively, while ensuring that the unit modulus constraints on $\bm{w}^{(t)}$, $\bm{v}_{\ell}^{(t+1)}$’s and the power constraints on $\{\bm{p}^{(t)},\bm{\lambda}^{(t)}\}$ are satisfied.
@@ V-A Simulation Setup
We consider an RIS-assisted MU-MISO communication system as illustrated in Fig. 8, consisting of an AP with $8$ antennas and an RIS with $100$ passive elements. In the Cartesian coordinate system shown in Fig. 8, the $(x,y,z)$-coordinates of the AP and the RIS are $(100\text{m},-100\text{m},0\text{m})$ and $(0\text{m},0\text{m},0\text{m})$, respectively. We assume that the RIS is equipped with a $10\times 10$ uniform rectangular array placed on the $(y,z)$-plane, and the AP has a uniform linear array parallel to the $x$-axis. The system is assumed to operate at a carrier frequency $f_{c}=1\text{GHz}$ with $10$MHz bandwidth. The uplink pilot transmit power $P_{u}$ and the downlink data transmit power $P_{d}$ are set to be $5$dBm and $10$dBm, and the noise spectrum density at the AP and UE are set to be $N_{0,u}=-154$dBm/Hz and $N_{0,d}=-160$dBm/Hz, respectively.
@@ V-B Implementation Details
The proposed active sensing framework is implemented on Tensorflow [40]. For the LSTM cells, the dimensions of hidden state $\bm{h}_{k}^{(t)}$’s, cell state $\bm{c}_{k}^{(t)}$’s, and hidden layers $\bm{u}_{c}$, $\bm{u}_{q}$, $\bm{u}_{i}$, $\bm{u}_{o}$, $\bm{w}_{c}$, $\bm{w}_{q}$, $\bm{w}_{i}$, $\bm{w}_{o}$ are set to be $512$. We adopt a GNN with $D=2$ updating layers. Within the GNN architecture, the DNNs $f_{1}(\cdot)$, $f_{2}(\cdot)$, and $f_{3}(\cdot)$ are of size $512\times 1024\times 512$, and the size of DNNs $f_{4}(\cdot)$, $f_{5}(\cdot)$, $f_{6}(\cdot)$, $f_{7}(\cdot)$, $f_{8}(\cdot)$, $f_{9}(\cdot)$, $f_{w}(\cdot)$, $f_{V}(\cdot)$ are $512\times 512\times 512$. We concatenate $U=20$ active sensing units to train the proposed framework. In each training epoch, the neural network samples $6400$ training data including random channel and noise realizations. We terminate the training process if the validation loss does not decrease over $30$ consecutive training epochs. An instance of trajectories for $K=3$ mobile UEs is shown in Fig. 8.
We assume there are $N_{G}=8$ paths between the AP and RIS. Each transmission frame contains $N=3$ blocks. In the pilot stage (the 00-th block) of each frame, there are $4$ sub-blocks, i.e., $L=3$. The effective downlink channel $\bm{h}_{c,k}^{(t,0)}$’s are estimated in the $4$-th sub-block with $\tau_{w}=10$ pilots.
@@ V-C Benchmark Methods
Figure 9: Average minimum user rate at the $1$-st block of each frame (i.e., $\min_{k}\{R_{k}^{(t,1)}\}$) for different beam tracking methods. Here, $L=3$, $M=8$, $N_{r}=100$, $K=3$, $N_{G}=8$, $\tau_{w}=10$.
Figure 10: The impact of different number of sensing vectors $L$’s within each frame on different beam tracking methods. Here, $M=8$, $N_{r}=100$, $K=3$, $N_{G}=8$, $\tau_{w}=10$.
@@ V-D Performance Comparison
Fig. 9 shows the downlink minimum user rate of different beam tracking methods in different transmission frames. The proposed active sensing framework initially achieves around $84\%$ of the BCD algorithm with perfect CSI, indicating efficient initial design based on the statistics of the UE mobility pattern. As the AP receives more pilots, the performance of the proposed approach improves quickly to $94\%$ of the performance of the benchmark with perfect CSI. This indicates that the proposed active sensing framework can efficiently exploit the temporal correlations of the time varying channel. Specifically, the beam tracking methods using both GNN and LSTM significantly outperform those benchmarks with only GNNs. This is because the LSTM-based methods are able to exploit the historical channel information to design the AP beamforming vectors and the downlink RIS reflection coefficients, thereby improving beam tracking performance over time. Among the beam tracking approaches with LSTM, the proposed framework performs the best, because it designs the RIS sensing vectors in an active manner. This shows the benefits of adaptively designing the RIS sensing vectors based on previously received pilots.
For the benchmarks with fixed sensing schemes, Fig. 9 shows that the ones with learned sensing vectors outperform those with random sensing vectors. This shows the performance gain of designing the RIS sensing vectors based on channel statistics. Additionally, the GNN & LSTM benchmark, which uses the same RIS reflection coefficients for both pilot acquisition and data transmission stages, performs less effectively as compared to other GNN & LSTM benchmarks that employ distinct RIS reflection coefficients for these two stages. This shows the importance of designing the RIS reflection coefficients in channel sensing and data transmission stages separately. Moreover, we remark that a $5\%$ performance improvement can be obtained from power allocation refinement.
Furthermore, Fig. 9 shows that the proposed framework maintains its performance for transmission frames beyond $20$, despite being trained with only $20$ concatenated active sensing units corresponding to $20$ frames. We also observe empirically that the proposed framework continues to sustain its performance over hundreds of frames by reusing the active sensing unit. This robust performance beyond the training range is due to the ability of the LSTM cells to effectively capture the UE mobility patterns and to consistently update channel information based on the latest received pilots.
@@
Next, we evaluate the performance of the proposed approach for different number of sensing vector $L$’s within each frame. As can be seen from Fig. 10, the proposed active sensing framework always outperforms the benchmarks with fixed sensing strategies, because it leverages past observations to actively design the RIS sensing vectors. Furthermore, it is observed that LSTM-based methods (including the proposed approach) show faster convergence as the number of sensing sub-blocks $L$ increases. This is because a larger $L$ provides the LSTM with more observations in every frame, thereby facilitating more efficient capturing of temporal channel correlations.
@@ V-E Interpretation of the Learned Solutions
Interpretation of the Downlink Beam Alignment Strategy
In Fig. 11(a), the array response of the downlink RIS reflection coefficients $\bm{w}^{(t)}$ shows that the proposed framework effectively tracks the UEs by progressively reflecting increasingly focused downlink beams via the RIS towards the actual positions of the mobile UEs. It is observed that the peaks corresponding to the first two UEs are sometimes weaker than the third UE, but we can see from Fig. 11(b) that these weaker UEs are compensated by stronger array response of the beamformers at the AP. Therefore, the proposed framework indeed learns to jointly optimize the downlink RIS reflection coefficients and the beamforming matrix to ensure fairness in maximizing the minimum rate across the three mobile UEs, while exploiting channel correlations to design improved RIS reflection coefficients and beamformers over time.
Interpretation of the Learned Solutions via SINR Map
@@
As illustrated in Fig. 13, the proposed framework designs $L=3$ RIS sensing vectors cooperatively. For example, in the pilot stage of the $30$-th frame, one UE (represented by the red dot) is not targeted by the first RIS sensing vector $\bm{v}_{1}^{(t)}$, but is focused by $\bm{v}_{2}^{(t)}$. Meanwhile, the weaker array response of $\bm{v}_{2}^{(t)}$ towards the green UE and $\bm{v}_{3}^{(t)}$ towards the blue UE are compensated by stronger responses from the other two sensing vectors, respectively. This cooperative sensing strategy is not seen in the initial tracking frame, but is progressively developed over time by leveraging historical channel observations. As a result, the proposed framework actively narrows the sensing range and enhances the pilot SNR in the uplink CSI acquisition stages, leading to more informative neural network inputs and improved beam tracking performance. In contrast, the GNN & LSTM benchmark employs a non-adaptive approach by using channel statistics to train fixed RIS sensing vectors, which cover nearly the entire service area for CSI acquisition. However, this broad coverage results in lower uplink SNR during CSI acquisition and, subsequently, higher pilot overhead.
VI Conclusion
(a) Array response of the downlink RIS reflection coefficients $\bm{w}^{(t)}$ obtained by the proposed framework.
(b) Array response of the downlink beamforming vectors $\bm{b}^{(t)}_{k}$’s obtained by the proposed framework.
Figure 11: Array response of the downlink RIS reflection coefficients and beamformers obtained by the proposed framework in different frames. The proposed framework is tested on a randomly generated trajectory. Here, $L=3$, $M=8$, $N_{r}=100$, $K=3$, $N_{G}=8$, $\tau_{w}=10$.